# Gravitational wave probes of particle dark matter: a review

### v4

@@ X.2 Constraints from sub-solar mass binary black hole searches
The lack of smoking-gun evidence for dark BHs in current LIGO, Virgo and KAGRA observations motivates the need to consider sub-solar mass BHs, whose origins, if detected, cannot be explained by astrophysical means. Searches for GWs from sub-solar mass compact objects in binary systems have grown in popularity in recent years, covering mass ranges of roughly $10^{-11},1]M_{\odot}$ Nitz and Wang (2021); Phukon et al. (2021); Abbott et al. (2022f); Miller et al. (2022a); Nitz and Wang (2022); Abbott et al. (2023, 2022b); Miller et al. (2024), and, traditionally, have put constraints on the fraction of DM that PBHs could compose. However, as discussed in Singh et al. (2021), these searches do not assume a particular formation mechanism for these compact objects, and the binaries can thus be interpreted as arising from not only PBHs, but also dark BHs.
@@
From this work Abbott et al. (2023), the mass of the dark proton has been constrained to be [0.66,8.8] GeV, which is actually broader than the constraint in Singh et al. (2021) of [0.91,4.76] GeV since there are no assumed detections of dark BHs.
@@
where $\alpha_{\chi}$ is the dark fine-structure constant and $m_{e_{\chi}}$ the mass of the dark electron. By combining the Chandrasekhar mass relation with assumptions about the dark fine-structure constant, the dark fermion masses, and the Jeans mass for collapse, one can use Eq. (1) of Singh et al. (2021) together with Eq. (4) of Shandera et al. (2018) to infer the allowed range of $\Delta E$. For the GW190425 interpretation, this yields a characteristic value of order $\sim 10^{-3}\,\mathrm{eV}$.
Figure 28: Taken from Abbott et al. (2023). Constraints on the fraction of DM that dark BHs could compose in the atomic DM scenario. A mass distribution for dark BHs is assumed with the shown minimum mass on the $x$-axis, and the constraints are shown for three different matched filtering algorithms: GstLAL (dotted), MBTA (dashed) and PyCBC (solid). Constraints from a previous search for sub-solar mass compact objects in the first half of the third observing run of LIGO, Virgo and KAGRA are also shown  Abbott et al. (2022f). This plot differs from Fig. 27 primarily in the mass range that is considered in the sub-solar mass search versus the ones in, e.g. Abbott et al. (2021c). At lower masses, the constraints on $f_{\rm DBH}$ weaken with respect to those at higher masses.
Although sub-solar mass searches for dark BHs have placed promising constraints, they are typically limited to masses above $\sim 10^{-2}M_{\odot}$, largely due to the prohibitive computational cost of conducting matched filtering searches for lighter-mass binaries. In contrast, CW methods—which search for time-frequency tracks rather than relying on matched filtering —can access lower-mass regimes Miller et al. (2021b); Abbott et al. (2022b); Miller (2024b); Alestas et al. (2024). While originally developed for PBH searches, these techniques are equally applicable to testing the dark BH hypothesis. As such, CW approaches offer a valuable path toward exploring a broader parameter space than is feasible with traditional matched filtering methods.
@@ XI Conclusions
Despite the “WIMP miracle” announced decades ago, WIMPs have not yet been detected, and are only one of numerous possible ways that DM could exist in the Universe. The lack of detection of WIMPs over the years has opened up both the experimental and theoretical playgrounds for new probes of DM, one of which is through its impact on GW signals, or as a signal in interferometer data itself. Moreover, it has been assumed in this work, and almost all those cited, that DM is the way to solve the puzzling cosmological observations on a variety of scales. However, modified theories of gravity are real contenders to solve this problem as well. These two communities – DM and modified gravity – act entirely independently, choosing to break, respectively, one of the two assumptions of “classical” physics: (1) the matter in the universe is mostly baryonic and luminous, and (2) general relativity and, in the limit of slow-moving objects, Newtonian theory, are correct Martens and Lehmkuhl (2020). It may be that these assumptions need to be broken simultaneously, or that these two distinct fields have more in common than it appears, as argued in Martens and Lehmkuhl (2020).
We only briefly mentioned PBHs in this review, which have been reviewed extensively elsewhere Green and Kavanagh (2021); Miller (2024a). PBHs could also comprise a fraction of or the totality of DM, and recent observations of binary black holes by the LIGO, Virgo and KAGRA with low spins whose merging rates are consistent with PBHs have reignited interested in this hypothesis Bird et al. (2016); Clesse and García-Bellido (2017); Sasaki et al. (2016); Croon et al. (2023). While analyses of PBHs typically receive criticism for using monochromatic mass functions to place constraints, all of the results presented here implicitly assume that DM is only composed of particles that are being constrained [^13] The exception to this statement is superradiance probes of BHs– these bosons need not be DM.. It could be that multiple forms of DM co-exist, e.g. in the atomic DM models, or that there is a whole dark sector.

### v5

@@ X.2 Constraints from sub-solar mass binary black hole searches
The lack of smoking-gun evidence for dark BHs in current LIGO, Virgo and KAGRA observations motivates the need to consider sub-solar mass BHs, whose origins, if detected, cannot be explained by astrophysical means. Searches for GWs from sub-solar mass compact objects in binary systems have grown in popularity in recent years, covering mass ranges of roughly $10^{-11},1]M_{\odot}$ Nitz:2021mzz; Phukon:2021cus; LIGOScientific:2021job; Miller:2021knj; Nitz:2022ltl; LIGOScientific:2022hai; KAGRA:2022dwb; Miller:2024fpo, and, traditionally, have put constraints on the fraction of DM that PBHs could compose. However, as discussed in Singh:2020wiq, these searches do not assume a particular formation mechanism for these compact objects, and the binaries can thus be interpreted as arising from not only PBHs, but also dark BHs.
@@
From this work LIGOScientific:2022hai, the mass of the dark proton has been constrained to be [0.66,8.8] GeV, which is actually broader than the constraint in Singh:2020wiq of [0.91,4.76] GeV since there are no assumed detections of dark BHs.
@@
where $\alpha_{\chi}$ is the dark fine-structure constant and $m_{e_{\chi}}$ the mass of the dark electron. By combining the Chandrasekhar mass relation with assumptions about the dark fine-structure constant, the dark fermion masses, and the Jeans mass for collapse, one can use Eq. (1) of Singh:2020wiq together with Eq. (4) of Shandera:2018xkn to infer the allowed range of $\Delta E$. For the GW190425 interpretation, this yields a characteristic value of order $\sim 10^{-3}\,\mathrm{eV}$.
Figure 28: Taken from LIGOScientific:2022hai. Constraints on the fraction of DM that dark BHs could compose in the atomic DM scenario. A mass distribution for dark BHs is assumed with the shown minimum mass on the $x$-axis, and the constraints are shown for three different matched filtering algorithms: GstLAL (dotted), MBTA (dashed) and PyCBC (solid). Constraints from a previous search for sub-solar mass compact objects in the first half of the third observing run of LIGO, Virgo and KAGRA are also shown  LIGOScientific:2021job. This plot differs from Fig. 27 primarily in the mass range that is considered in the sub-solar mass search versus the ones in, e.g. LIGOScientific:2021djp. At lower masses, the constraints on $f_{\rm DBH}$ weaken with respect to those at higher masses.
Although sub-solar mass searches for dark BHs have placed promising constraints, they are typically limited to masses above $\sim 10^{-2}M_{\odot}$, largely due to the prohibitive computational cost of conducting matched filtering searches for lighter-mass binaries. In contrast, CW methods—which search for time-frequency tracks rather than relying on matched filtering —can access lower-mass regimes Miller:2020kmv; KAGRA:2022dwb; Miller:2024khl; Alestas:2024ubs. While originally developed for PBH searches, these techniques are equally applicable to testing the dark BH hypothesis. As such, CW approaches offer a valuable path toward exploring a broader parameter space than is feasible with traditional matched filtering methods.
@@ XI Conclusions
Despite the “WIMP miracle” announced decades ago, WIMPs have not yet been detected, and are only one of numerous possible ways that DM could exist in the Universe. The lack of detection of WIMPs over the years has opened up both the experimental and theoretical playgrounds for new probes of DM, one of which is through its impact on GW signals, or as a signal in interferometer data itself. Moreover, it has been assumed in this work, and almost all those cited, that DM is the way to solve the puzzling cosmological observations on a variety of scales. However, modified theories of gravity are real contenders to solve this problem as well. These two communities – DM and modified gravity – act entirely independently, choosing to break, respectively, one of the two assumptions of “classical” physics: (1) the matter in the universe is mostly baryonic and luminous, and (2) general relativity and, in the limit of slow-moving objects, Newtonian theory, are correct Martens:2020lto. It may be that these assumptions need to be broken simultaneously, or that these two distinct fields have more in common than it appears, as argued in Martens:2020lto.
We only briefly mentioned PBHs in this review, which have been reviewed extensively elsewhere Green:2020jor; Miller:2024rca. PBHs could also comprise a fraction of or the totality of DM, and recent observations of binary black holes by the LIGO, Virgo and KAGRA with low spins whose merging rates are consistent with PBHs have reignited interested in this hypothesis Bird:2016dcv; Clesse:2016vqa; Sasaki:2016jop; Croon:2022tmr. While analyses of PBHs typically receive criticism for using monochromatic mass functions to place constraints, all of the results presented here implicitly assume that DM is only composed of particles that are being constrained [^13] The exception to this statement is superradiance probes of BHs– these bosons need not be DM.. It could be that multiple forms of DM co-exist, e.g. in the atomic DM models, or that there is a whole dark sector.